# Meta-Transfer Derm-Diagnosis: Exploring Few-Shot Learning and Transfer Learning for Skin Disease Classification in Long-Tail Distribution

### v1

@@ 6 Conclusion and Future Work
In this study, we explore the effectiveness of episodic and traditional training methods combined with few-shot learning through transfer learning for classifying rare skin diseases in situations with limited data. Current literature emphasize that datasets with long-tail distributions present specific challenges, highlighting the necessity for specialized training approaches to achieve successful outcomes. But there is a lack of research combining basic few-shot training models with well-known transfer learning methodologies in this domain. Our study seeks to fill this gap by evaluating these established training methods on major datasets often used as benchmarks for long-tail distributions in skin diseases. In this context, we presented evaluations of three publicly available skin image datasets using the FETL, FEL, DTL, and DL models. Notably, the DTL model, which integrates transfer learning with conventional training techniques, consistently outperformed others in these datasets. By enhancing our models with MixUp, CutMix, and ResizeMix techniques, as recommended in existing literature, we were able to reach leading performance on the SD-198 and comparable performances with the state-of-the-art methods in ISIC2018 and Derm7pt datasets.
These findings underline the effectiveness of transfer learning, a well-established method, in the context of few-shot learning for long-tail distributions in skin disease classification. We suggest that further improvements in accuracy and robustness can be achieved through expanded data augmentation and exploring different model architectures. Based on the insights gained from this study, we believe that further research is crucial, especially for dermoscopic datasets facing data scarcity. Moreover, applying these findings to a broader range of medical imaging datasets, including X-rays, CT scans, and MRIs, could be helpful in detecting and diagnosing new diseases.

### v2

@@ 7 Conclusion and Future Work
In this study, we explored the effectiveness of supervised and self-supervised transfer learning strategies combined with few-shot learning to classify rare skin diseases under long-tail data distributions. While prior work has highlighted the challenges of imbalance and data scarcity in this domain, there remains limited research evaluating the combined impact of transfer learning and few-shot methods in a unified setting. To address this gap, we systematically evaluated five training paradigms (FETL, FEL, DTL, DTLS, and DL) across three benchmark skin image datasets. Our findings demonstrate that DTL models, particularly those using MobileNetV2 backbones and enhanced with MixUp, CutMix, and ResizeMix, consistently outperformed alternative approaches. In addition, transformer-based models such as ViT-Base, fine-tuned via LoRA, achieved state-of-the-art performance on the Derm7pt dataset.
These results reaffirm the strength of conventional transfer learning in few-shot medical imaging tasks and highlight the value of appropriate data augmentation and backbone selection. While episodic and self-supervised methods showed advantages in specific low-shot scenarios, their performance was generally less stable compared to supervised DTL approaches. Additionally, contrastive self-supervised pretraining strategies such as SimCLR and MoCo-v3 were also applied and yielded competitive results in selected scenarios. The datasets used in this study only contained class labels and did not provide additional clinical context such as lesion descriptions, anatomical site annotations, or physician notes. For this reason, vision–language foundation models (e.g., BioViL, MedCLIP) were evaluated in a zero-shot setting and showed limited transferability due to domain mismatch and the absence of text-based adaptation.
Future research may benefit from exploring domain-specific self-supervised pretraining strategies, lightweight transformer fine-tuning methods, and the integration of multimodal approaches. Recently introduced dermatology datasets with rich textual annotations, such as DermaCon-IN [83], indicate that the field is moving toward cross-modal resources. Building upon these, future work could explore integration with CLIP-based vision–language models and multimodal LLMs such as GPT-4o or BiomedCLIP to combine lesion images with textual or patient-level context, enabling interactive diagnostic support and zero-shot generalization. Finally, extending the framework to other medical imaging modalities (e.g., X-rays, histopathology, fundus imaging) in parallel with multimodal integration could further support robust classification systems for low-data and rare disease applications.
Author Biographies
Zeynep Özdemir received her B.S. and M.S. degrees in Computer Engineering from Ankara University, Turkey, in 2016 and 2018, respectively, and is currently pursuing her Ph.D. in Computer Engineering at the same institution. Since 2018, she has been a Research Assistant in the Department of Computer Engineering at Ankara University. Her research interests lie in artificial intelligence for healthcare, with a focus on computer vision, medical image analysis, explainable AI, segmentation, domain adaptation, and rare disease classification. Her doctoral research particularly explores few-shot and transfer learning under long-tail distributions for skin disease classification. She actively engages with the international MICCAI community through the RISE-MICCAI reading group, further expanding her expertise in medical imaging and computational methods.
Dr. Hacer Yalım Keleş completed her B.Sc., M.Sc., and Ph.D. in Computer Engineering at Middle East Technical University, Turkey, in 2002, 2005, and 2010, respectively. Her Ph.D. thesis was honored with the Thesis of the Year award by the Prof. Dr. Mustafa Parlar Education and Research Foundation in 2010. Between 2000 and 2007, she contributed as a researcher at The Scientific and Technological Research Council of Turkey (TUBITAK), focusing on pattern recognition using multimedia data, including audio and video.
Dr. Keles was an Assistant Professor at the Department of Computer Engineering, Ankara University, from 2013 to 2021, and is currently an Associate Professor at the Department of Computer Engineering, Hacettepe University.
Her research primarily spans computer vision and machine learning, with a focus on learning algorithms for limited data and deep generative models. She has contributed to sign and gesture recognition, generative adversarial networks, image inpainting, and image segmentation domains. Moreover, she collaborates on diverse projects involving aerial and medical images, speech signals, textual, geophysical, and hyperspectral data analysis with her graduate students.
Dr. Ömer Özgür Tanrıöver received the B.Sc. degree in computer engineering, the M.Sc. and Ph.D. degrees in information systems from Middle East Technical University, Ankara, Turkey. Previously, he has served as a Certified Information Systems Auditor (CISA) with the Information Management Department, Banking Regulation Agency of Turkey. Currently, he is with Computer Engineering Department of Ankara University as an Associate Professor. His current research interests include applications of machine learning, information systems security and software process.